# Multi-view Structural Convolution Network for Domain-Invariant Point Cloud Recognition of Autonomous Vehicles

### v2

@@ Abstract
I Introduction
@@
Moreover, ensuring the robust and consistent trained model is essential for perception, localization and mapping in autonomous vehicles. As depicted in Fig 1, the quality of point clouds obtained through LiDAR sensors heavily depends on the sensor’s resolution. However, LiDAR sensors vary significantly in resolution depending on the manufacturer and its model, leading to substantial differences in the quality of point clouds used for recognition models. Furthermore, real-world objects exhibit significant variation based on its installed location or multi-sensor usage, adding another layer of complexity. These disparities in real-world point clouds have different point cloud patterns and resolutions and result in misjudgments on perception in self-driving vehicles. For example, during the development phase, self-driving vehicles may be trained using high-resolution LiDAR sensors to produce high-quality point clouds. However, in practical deployment scenarios, commercial vehicles may be equipped with low-resolution LiDAR sensors due to cost or other constraints. This disparity can lead to severe performance gaps between the development and deployment environments. Furthermore, when self-driving vehicles are deployed in new unvisited regions or environments during training, models can experience a significant drop in performance[30].
@@
II related works
II-A 3D point cloud deep learning networks: Encoding perspectives
II-A1 Multi-view-based methods
@@
II-A2 Voxel-based methods
@@
II-A3 Point cloud-based methods
@@
II-B 3D point cloud deep learning networks: Local aggregation
II-B1 Point-by-point methods
@@
II-B2 Graph-based methods
@@
II-C Domain adaptation and Domain generalization algorithms
II-C1 Domain adaptation
@@
II-C2 Single domain generalization
@@
III Multi-view Strucutral Convolution Network(MSCN)
@@
III-A Structural Convolution Layer(SCL)
@@
$Conv_{dir}(R^{M}_{n},K^{S}_{dir})=\langle F(p_{n}),w(b^{c}_{dir})\rangle+\sum\limits_{s=1}^{S}\max\limits_{m\in(1,M)}sim(p_{m},b^{s}_{dir})$
@@
$SCL(D_{i-1},D_{i})=\\ MLP(Conv_{dir}(R^{M}_{n},K^{S}_{dir})\oplus Conv_{dist}(R^{M}_{n},K^{S}_{dist}))$
III-B Structural Aggregation Layer (SAL)
@@ III-C Classification Process
IV PROGRESSIVE DOMAIN EXPANSION PROCESS
@@ IV-A Generator composition
IV-B MSCN training with $\{S^{{}^{\prime}}\}^{k-1}_{i=1}$
@@
$L_{ce}(x_{i},y_{i})$ $=\min\limits_{F,C}-\sum_{m=1}^{M}y^{m}_{i}log(C(F(x^{m}_{i})))$ (12)
$L_{nce}(z_{i},z^{+}_{i})$ $=\min\limits_{F,C}-log\frac{exp(z_{i}\cdot z^{+}_{i})}{\sum^{2N}_{j=1,j\neq i}exp(z_{i}\cdot z_{j})}$
@@
IV-C Unseen Domain $S^{{}^{\prime}}_{k}$ Generation
To effectively and safely generate domains, we compose the unseen domain generation process as shown in Figure 3 (b). The source domain point cloud $x$ is converted to an unseen domain point cloud $x^{{}^{\prime}}_{k}$ in the domain generation module $G_{k}$, which is then used as input for both the MSCN and the reconstruction module $G_{R}$. The primary goal of the reconstruction module is to reconstruct $x^{{}^{\prime}}_{k}$ to $x$i to ensure the safety of the decoder $G_{D}$ and $AdaIN_{3D}$ in $G_{k}$. This is achieved by minimizing reconstruction loss defined as:
@@ V EVALUATION
V-A Datasets & Implementation Detail
V-A1 Synthetic Datasets
@@
V-A2 Real-world Indoor Dataset
@@
V-A3 Real-World Self-Driving Vehicle Datasets
@@
Comparisons between our methods are conducted, including MSCN and MSCN, and precious methods including PointNet[1], PointNeXt[29], PointMLP[4], DGCNN[3], GCN[5], SCN[30], PointTransformer[36], and PointMamba[35]. PointMamba was first pre-trained on ShapeNet[37] and then trained on KITTI, PanKyo, and nuScenes for the classification task. The newtworks were evalauted through a standard object classification accuracy.
@@
TABLE I: Object part segmentation results on ShapeNetPart[21], grouped by dataset. mIOU values are reported for each method and dataset.
V-B Evaluation on Synthetic Datasets: Robustness to geometric deformation
We first conducted experiments on the ShapeNetPart [21] dataset for part segmentation. Table I summarizes the part segmentation results, where mIoU values are reported for each method. Our MSCN achieves performance comparable to state-of-the-art (SOTA) models, despite using only 1024 points, unlike other SOTA models that rely on 2048 points. While SOTA models exhibit high performance on the given dataset, MSCN’s performance is slightly inferior compared to SOTA models. This performance gap can be explained by the inherent trade-off between robustness and accuracy, as discussed in [57]. Our MSCN were designed to generalize well to harsh or out-of-distribution scenarios while preserving accuracy on natural data. However, Zhang et.al[57] showed this approach often leads to a shift in focus from minimizing standard classification errors (natural error) to improving robustness, which can degrade accuracy in ideal.
@@ V-B Evaluation on Synthetic Datasets: Robustness to geometric deformation
TABLE II: Classification results on ModelNet40 [27].
Furthermore, we conducted experiments on the ModelNet40 [27] dataset for object classification. Table II presents the classification results, where MSCN is slightly lower than some state-of-the-art (SOTA) models. Although SOTA models demonstrate strong performance on this dataset, their sensitivity to geometric transformations can result in significant performance degradation in more challenging scenarios. To further investigate this, Figure 5 illustrates the performance of models under geometric transformations such as translation and scaling. MSCN consistently exhibits robust performance across all transformations, maintaining stable accuracy despite severe geometric changes. In contrast, other methods experience substantial drops in accuracy, highlighting their vulnerability to such transformations. These results underscore the strength of MSCN in terms of geometrical invariance, demonstrating its potential for real-world applications where robustness to geometric variations is critical.
@@
TABLE III: Classification results on the ScanObjectNN[56] dataset. This table summarizes the performance of various methods under different experimental settings, including self-supervised pretraining, supervised learning, and the evaluation of MSCN’s translation and scale invariance. PT indicates Pretrained models, RT represents Random Translation, and RS denotes Random Scaling.
@@
TABLE IV: Intra-domain classification results on KITTI [25], nuScenes [26], and PanKyo [30] datasets. The bold black, blue indicate the first and the second best performance on each scenario, respectively.
@@
TABLE V: Cross-domain classification results on KITTI [25], nuScenes [26], PanKyo [30], and simulation [31] datasets. The bold black, blue indicate the first and the second best performance on each scenario, respectively.
@@
TABLE VI: Inference time per point cloud on KITTI [25], nuScenes [26], PanKyo [30], and simulation [31] datasets. All values are measured in milliseconds (ms).
@@
TABLE VII: Layer configuration evaluation results on KITTI [25], nuScenes [26], PanKyo [30], and simulation [31] datasets. The bold black indicates the best performance on each scenario.
@@
V-C Evaluation on Indoor Real-World Datasets: Comparable Performance and Robustness
We conducted experiments on the ScanObjectNN [56] dataset to evaluate the performance of MSCN on real-world indoor point clouds. Table III summarizes the classification results, demonstrating that MSCN achieves performance comparable to the state-of-the-art (SOTA) models in standard evaluation settings. Furthermore, MSCN exhibits remarkable robustness under geometric transformations such as scaling and translation, maintaining stable performance and demonstrating its strength in handling these variations.
V-D Evaluation on Self-Driving Datasets: Cross-Domain Robustness
First, the performances are demonstrated within the KITTI[25], nuScenes[26], and PanKyo[30] datasets to capture classification performance seamlessly using the training dataset on the test dataset. Table IV shows the classification performance within a single domain. These results measure the model’s ability to classify point clouds within the same domain used for training. While MSCN achieves competitive results, it lags slightly behind the SOTA methods in terms of accuracy. For example, the average intra-domain accuracy of MSCN across all datasets is 94.7%, compared to PointTransformer’s 98%. This performance gap can be explained from the bias-variance trade-off [38]. MSCN, designed for robust domain-invariant feature extraction, prioritizes reducing model variance to generalize effectively across diverse domains. As a result, its intra-domain performance exhibits slightly higher bias, which is a common trade-off observed in models designed with high generalization capacity. Specifically, while other SOTA models may minimize bias within the training domain, they often exhibit poor performance in domain-shift scenarios due to higher variance.
To demonstrate MSCN’s robustness, we further evaluated its performance under cross-domain scenarios as illustrated in Table V, where substantial domain shifts between training and testing datasets. PointNet, PointNeXt, PointMLP, DGCNN, PointTransformer, and PointMamba demonstrate good performance when the domain shifts from simulation to the real world (Sim2Real) but show significantly lower performance when shifting from one real-world dataset to another (Real2Real). On the other hand, GCN and SCN show robust performance in Real2Real domain change scenarios but are sensitive in Sim2Real scenarios.
@@ V-D Evaluation on Self-Driving Datasets: Cross-Domain Robustness
Furthermore, Table VI presents the inference time required to predict the class of a single point cloud in both intra-domain and cross-domain settings. Models such as PointMLP, PointNeXt, and PointTransformer exhibit higher inference times due to their significantly large number of parameters. In contrast, although PointMamba is also Transformer-based and has a large parameter count, its inference time is relatively low. This efficiency can be attributed to its design, which likely focuses on reducing FLOPs (Floating Point Operations Per Second) to enhance computational efficiency.
V-E Ablation study
Firstly, to assess the compatibility of layer configurations of MSCN, we compare the cases where there are configurations of both SCL and SAL, and only SCL. The former consists of four SCL and two SAL, which is illustrated in Figure 2 (a). The latter is composed of five SCL that are SCL(3, 32), SCL(32, 64), SCL(64, 128), SCL(128, 256), and SCL(256, 1024). Table VII shows the performance of the methods in layer configurations where domain shift scenarios. MSCN outperforms the case that has only SCL without SAL during all domain shift scenarios, with an average performance that is 7.2% higher than the case that only consist of SCL.
To evaluate the impact of our unseen domain generalization methods on improving the leaning of domain-invariant feature representation, we performed various experiments based on both intra-domain and cross-domain classification, which are illustrated in Table VIII. MSCN, which can increase the amount of training without overfitting, performs better than MSCN for single-domain classification. Note that MSCN does not include a domain generation module. With the domain generation module, MSCN demonstrates even more robust performance than MSCN in various domain change scenarios, with an average improvement of 5.8%, indicating that our proposed unseen domain generation method ensures domain generalization ability.
@@ Acknowledgment
This work was supported by the Korea National Police Agency(KNPA) under the project ’Development of the autonomous driving patrol service for active prevention and response to traffic accidents’ (RS-2024-00403630) and the BK21 FOUR program of the National Research Foundation Korea funded by the Ministry of Education(NRF5199991014091).

### v3

@@ Abstract
journal: Expert Systems with Applications
[…]
[a]organization=Department of Civil Engineering, University of Central Florida,postcode=32816, state=Florida, country=USA
[…]
[b]organization=Department of Data, Networks and AI, Ajou University,postcode=16499, state=Suwon, country=South Korea
[…]
[c]organization=Department of Mobility Engineering and Department of Artificial Intelligence, Ajou University,postcode=16499, state=Suwon, country=South Korea
1 Introduction
@@
Moreover, ensuring the robust and consistent trained model is essential for perception, localization and mapping in autonomous vehicles. As depicted in Figure 1, the quality of point clouds obtained through LiDAR sensors heavily depends on the sensor’s resolution. However, LiDAR sensors vary significantly in resolution depending on the manufacturer and its model, leading to substantial differences in the quality of point clouds used for recognition models. Furthermore, real-world objects exhibit significant variation based on its installed location or multi-sensor usage, adding another layer of complexity. These disparities in real-world point clouds have different point cloud patterns and resolutions and result in misjudgments on perception in self-driving vehicles. For example, during the development phase, self-driving vehicles may be trained using high-resolution LiDAR sensors to produce high-quality point clouds. However, in practical deployment scenarios, commercial vehicles may be equipped with low-resolution LiDAR sensors due to cost or other constraints. This disparity can lead to severe performance gaps between the development and deployment environments. Furthermore, when self-driving vehicles are deployed in new unvisited regions or environments during training, models can experience a significant drop in performance[30].
@@
2 related works
2.1 3D point cloud deep learning networks: Encoding perspectives
2.1.1 Multi-view-based methods
@@
2.1.2 Voxel-based methods
@@
2.1.3 Point cloud-based methods
@@
2.2 3D point cloud deep learning networks: Local aggregation
2.2.1 Point-by-point methods
@@
2.2.2 Graph-based methods
@@
2.3 Domain adaptation and Domain generalization algorithms
2.3.1 Domain adaptation
@@
2.3.2 Single domain generalization
@@
3 Multi-view Strucutral Convolution Network(MSCN)
@@
3.1 Structural Convolution Layer(SCL)
@@
$Conv_{dir}(R^{M}_{n},K^{S}_{dir})=\langle F(p_{n}),w(b^{c}_{dir})\rangle+\sum\limits_{s=1}^{S}\max\limits_{m\in(1,M)}sim(p_{m},b^{s}_{dir})$ (4)
@@
$SCL(D_{i-1},D_{i})=MLP(Conv_{dir}(R^{M}_{n},K^{S}_{dir})\oplus Conv_{dist}(R^{M}_{n},K^{S}_{dist}))$ (7)
3.2 Structural Aggregation Layer (SAL)
@@ 3.3 Classification Process
4 PROGRESSIVE DOMAIN EXPANSION PROCESS
@@ 4.1 Generator composition
4.2 MSCN training with $\{S^{{}^{\prime}}\}^{k-1}_{i=1}$
@@
$L_{ce}(x_{i},y_{i})$ $=\min\limits_{F,C}-\sum_{m=1}^{M}y^{m}_{i}\log(C(F(x^{m}_{i})))$ (12)
$L_{nce}(z_{i},z^{+}_{i})$ $=\min\limits_{F,C}-\log\left(\frac{\exp(z_{i}\cdot z^{+}_{i})}{\sum^{2N}_{j=1,j\neq i}\exp(z_{i}\cdot z_{j})}\right)$
@@
4.3 Unseen Domain $S^{{}^{\prime}}_{k}$ Generation
To effectively and safely generate domains, we compose the unseen domain generation process as shown in Figure 3 (b). The source domain point cloud $x$ is converted to an unseen domain point cloud $x^{{}^{\prime}}_{k}$ in the domain generation module $G_{k}$, which is then used as input for both the MSCN and the reconstruction module $G_{R}$. The primary goal of the reconstruction module is to reconstruct $x^{{}^{\prime}}_{k}$ to $x_{i}$ to ensure the safety of the decoder $G_{D}$ and $AdaIN_{3D}$ in $G_{k}$. This is achieved by minimizing reconstruction loss defined as:
@@ 5 EVALUATION
5.1 Datasets & Implementation Detail
5.1.1 Synthetic Datasets
@@
5.1.2 Real-world Indoor Dataset
@@
5.1.3 Real-World Self-Driving Vehicle Datasets
@@
Comparisons between our methods are conducted, including MSCN and MSCN, and precious methods including PointNet[1], PointNeXt[29], PointMLP[4], DGCNN[3], GCN[5], SCN[30], PointTransformer[36], and PointMamba[35]. PointMamba was first pre-trained on ShapeNet[37] and then trained on KITTI, PanKyo, and nuScenes for the classification task. The newtworks were evalauted through the a standard accuracy metric.
@@
Table 1: Object part segmentation results on ShapeNetPart[21], grouped by dataset. mIOU values are reported for each method and dataset.
5.2 Evaluation on Synthetic Datasets: Robustness to geometric deformation
We first conducted experiments on the ShapeNetPart [21] dataset for part segmentation. Table 1 summarizes the part segmentation results, where mIoU values are reported for each method. Our MSCN achieves performance comparable to state-of-the-art (SOTA) models, despite using only 1024 points, unlike other SOTA models that rely on 2048 points. While SOTA models exhibit high performance on the given dataset, MSCN’s performance is slightly inferior compared to SOTA models. This performance gap can be explained by the inherent trade-off between robustness and accuracy, as discussed in [57]. Our MSCN were designed to generalize well to harsh or out-of-distribution scenarios while preserving accuracy on natural data. However, Zhang et.al[57] showed this approach often leads to a shift in focus from minimizing standard classification errors (natural error) to improving robustness, which can degrade accuracy in ideal.
@@ 5.2 Evaluation on Synthetic Datasets: Robustness to geometric deformation
Table 2: Classification results on ModelNet40 [27].
Furthermore, we conducted experiments on the ModelNet40 [27] dataset for object classification. Table 2 presents the classification results, where MSCN is slightly lower than some state-of-the-art (SOTA) models. Although SOTA models demonstrate strong performance on this dataset, their sensitivity to geometric transformations can result in significant performance degradation in more challenging scenarios. To further investigate this, Figure 5 illustrates the performance of models under geometric transformations such as translation and scaling. MSCN consistently exhibits robust performance across all transformations, maintaining stable accuracy despite severe geometric changes. In contrast, other methods experience substantial drops in accuracy, highlighting their vulnerability to such transformations. These results underscore the strength of MSCN in terms of geometrical invariance, demonstrating its potential for real-world applications where robustness to geometric variations is critical.
@@
Table 3: Classification results on the ScanObjectNN[56] dataset. This table summarizes the performance of various methods under different experimental settings, including self-supervised pretraining, supervised learning, and the evaluation of MSCN’s translation and scale invariance. PT indicates Pretrained models, RT represents Random Translation, and RS denotes Random Scaling.
5.3 Evaluation on Indoor Real-World Datasets: Comparable Performance and Robustness
We conducted experiments on the ScanObjectNN [56] dataset to evaluate the performance of MSCN on real-world indoor point clouds. Table 3 summarizes the classification results, demonstrating that MSCN achieves performance comparable to the state-of-the-art (SOTA) models in standard evaluation settings. Furthermore, MSCN exhibits remarkable robustness under geometric transformations such as scaling and translation, maintaining stable performance and demonstrating its strength in handling these variations.
@@ 5.3 Evaluation on Indoor Real-World Datasets: Comparable Performance and Robustness
Table 4: Intra-domain classification results on KITTI [25], nuScenes [26], and PanKyo [30] datasets. The bold black, blue indicate the first and the second best performance on each scenario, respectively.
5.4 Evaluation on Self-Driving Datasets: Cross-Domain Robustness
First, the performances are demonstrated within the KITTI[25], nuScenes[26], and PanKyo[30] datasets to capture classification performance seamlessly using the training dataset on the test dataset. Table 4 shows the classification performance within a single domain. These results measure the model’s ability to classify point clouds within the same domain used for training. While MSCN achieves competitive results, it lags slightly behind the SOTA methods in terms of accuracy. For example, the average intra-domain accuracy of MSCN across all datasets is 94.7%, compared to PointTransformer’s 98%. This performance gap can be explained from the bias-variance trade-off [38]. MSCN, designed for robust domain-invariant feature extraction, prioritizes reducing model variance to generalize effectively across diverse domains. As a result, its intra-domain performance exhibits slightly higher bias, which is a common trade-off observed in models designed with high generalization capacity. Specifically, while other SOTA models may minimize bias within the training domain, they often exhibit poor performance in domain-shift scenarios due to higher variance.
@@ 5.4 Evaluation on Self-Driving Datasets: Cross-Domain Robustness
Table 5: Cross-domain classification results on KITTI [25], nuScenes [26], PanKyo [30], and simulation [31] datasets. The bold black, blue indicate the first and the second best performance on each scenario, respectively.
@@
Table 6: Inference time per point cloud on KITTI [25], nuScenes [26], PanKyo [30], and simulation [31] datasets. All values are measured in milliseconds (ms).
To demonstrate MSCN’s robustness, we further evaluated its performance under cross-domain scenarios as illustrated in Table 5, where substantial domain shifts between training and testing datasets. PointNet, PointNeXt, PointMLP, DGCNN, PointTransformer, and PointMamba demonstrate good performance when the domain shifts from simulation to the real world (Sim2Real) but show significantly lower performance when shifting from one real-world dataset to another (Real2Real). On the other hand, GCN and SCN show robust performance in Real2Real domain change scenarios but are sensitive in Sim2Real scenarios.
Furthermore, Table 6 presents the inference time required to predict the class of a single point cloud in both intra-domain and cross-domain settings. Models such as PointMLP, PointNeXt, and PointTransformer exhibit higher inference times due to their significantly large number of parameters. In contrast, although PointMamba is also Transformer-based and has a large parameter count, its inference time is relatively low. This efficiency can be attributed to its design, which likely focuses on reducing FLOPs (Floating Point Operations Per Second) to enhance computational efficiency.
@@
Table 7: Layer configuration evaluation results on KITTI [25], nuScenes [26], PanKyo [30], and simulation [31] datasets. The bold black indicates the best performance on each scenario.
5.5 Ablation study
Firstly, to assess the compatibility of layer configurations of MSCN, we compare the cases where there are configurations of both SCL and SAL, and only SCL. The former consists of four SCL and two SAL, which is illustrated in Figure 2 (a). The latter is composed of five SCL that are SCL(3, 32), SCL(32, 64), SCL(64, 128), SCL(128, 256), and SCL(256, 1024). Table 7 shows the performance of the methods in layer configurations where domain shift scenarios. MSCN outperforms the case that has only SCL without SAL during all domain shift scenarios, with an average performance that is 7.2% higher than the case that only consist of SCL.
@@ 5.5 Ablation study
To evaluate the impact of our unseen domain generalization methods on improving the leaning of domain-invariant feature representation, we performed various experiments based on both intra-domain and cross-domain classification, which are illustrated in Table 8. MSCN, which can increase the amount of training without overfitting, performs better than MSCN for single-domain classification. Note that MSCN does not include a domain generation module. With the domain generation module, MSCN demonstrates even more robust performance than MSCN in various domain change scenarios, with an average improvement of 5.8%, indicating that our proposed unseen domain generation method ensures domain generalization ability.
@@ Acknowledgment
This work was supported by the Korea National Police Agency(KNPA) under the project ’Development of the autonomous driving patrol service for active prevention and response to traffic accidents’ (RS-2024-00403630).